# Márcia Lemos-Silva1

### v1

@@ 1 Introduction
where $b,c\in\mathbb{R}^{+}$, and where $x(t_{0})=x_{0}$, $y(t_{0})=y_{0}$, and $z(t_{0})=z_{0}$ are nonnegative and, moreover, $x,y,z:\mathbb{R}\rightarrow\mathbb{R}^{+}_{0}$. This model adopts a well-known compartmental approach, wherein three variables represent different groups within the population. The variable $x$ represents the proportion of susceptible individuals, $y$ represents the proportion of infected individuals, and $z$ represents the proportion of removed individuals. By adding the right-hand side of equations (1), one has $x^{\prime}(t)+y^{\prime}(t)+z^{\prime}(t)=0$, which is equivalent to $x(t)+y(t)+z(t)=N$, for all $t$, where the constant $N:=x_{0}+y_{0}+z_{0}$ denotes the total population under study. The exact solution of (1) was found in [5]. More recently, a different method was proposed to solve (1), generalizing the result of [5] to non-autonomous cases [2].
@@
Here we propose and investigate a SIR model formulated in quantum time. One of the main motivations for considering such a model lies in the non-uniform structure of the quantum time scale. Unlike the standard discrete-time framework, where time progresses in uniform steps, quantum time evolves on a increasing step. This means that during the early stages of an epidemic, where dynamics tend to be more sensitive, the step is finer, turning to a coarser one as the system stabilizes. Quantum modeling offers a more natural discretization for many real-world phenomena, where the initial dynamics changes rapidly before approaching an equilibrium. To the best of our knowledge, no SIR model has yet been formulated in quantum time, making our approach a novel contribution to the field. Thus, regarding applications, the quantization of infectious disease modeling may effectively address some limitations of traditional continuous compartmental models by introducing a novel approach to understand disease dynamics. Furthermore, unlike many SIR models, our proposed model has an exact solution. Finally, we prove that all its solutions remain non-negative and that their long-term behavior are coherent with (1), which guarantees consistency.
@@ 2.3 Exact solution
Here we derive the exact solution of system (4). Note that, since $N=x(qt)+y(qt)+z(qt)$ for all $t$, it means it is enough to solve the first two equations of system (4): after we know $x(t)$ and $y(t)$, we can immediately compute $z(qt)$ using the equality z(qt)=Nx(qt)y(qt)z(qt)=N-x(qt)-y(qt).
@@
which tends to zero as $i$ grows. Now we prove that $\xi_{i+1}<\xi_{i}<1$ for all $i\geq 0$. For a certain $i\geq 0$, suppose that $\xi_{i+1}<\xi_{i}$. Then,
@@
Since, by hypothesis, $\mathcal{R}_{0}<1$, then $\frac{c}{b}>1$, and thus the result holds for any $q>1$. ∎
@@ 4 Conclusion
In this work we proposed a quantum SIR model considering the concept of $q$-derivative and a particular rule of Mickens’ method. The equilibrium points were found and the non-negativity and boundedness of solutions were proved. We derived the exact solution of the quantum SIR model and proved asymptotic stability. We end providing some examples that illustrate the exact solution obtained.
@@
The authors declare no conflict of interest.

### v2

@@ 1 Introduction
where $b,c\in\mathbb{R}^{+}$, and where $x(t_{0})=x_{0}$, $y(t_{0})=y_{0}$, and $z(t_{0})=z_{0}$ are non-negative and, moreover, $x,y,z:\mathbb{R}\rightarrow\mathbb{R}^{+}_{0}$. This model adopts a well-known compartmental approach, wherein three variables represent different groups within the population. The variable $x$ represents the proportion of susceptible individuals, $y$ represents the proportion of infected individuals, and $z$ represents the proportion of removed individuals. By adding the right-hand side of equations (1), one has $x^{\prime}(t)+y^{\prime}(t)+z^{\prime}(t)=0$, which is equivalent to $x(t)+y(t)+z(t)=N$, for all $t$, where the constant $N:=x_{0}+y_{0}+z_{0}$ denotes the total population under study. The exact solution of (1) was found in [5]. More recently, a different method was proposed to solve (1), generalizing the result of [5] to non-autonomous cases [2].
@@
Here, we propose and investigate a SIR model formulated in quantum time. One of the main motivations for considering such a model lies in the non-uniform structure of the quantum time scale. Unlike the standard discrete-time framework, where time progresses in uniform steps, quantum time evolves on a increasing step. This means that during the early stages of an epidemic, where dynamics tend to be more sensitive, the step is finer, turning to a coarser one as the system stabilizes. Quantum modeling offers a more natural discretization for many real-world phenomena, where the initial dynamics changes rapidly before approaching an equilibrium. To the best of our knowledge, no SIR model has yet been formulated in quantum time, making our approach a novel contribution to the field. Thus, regarding applications, the quantization of infectious disease modeling may effectively address some limitations of traditional continuous compartmental models by introducing a novel approach to understand disease dynamics. Furthermore, unlike many SIR models, our proposed model has an exact solution. Finally, we prove that all its solutions remain non-negative and that their long-term behavior are coherent with (1), which guarantees consistency.
@@ 2.3 Exact solution
Here, we derive the exact solution of system (4). Note that, since $N=x(qt)+y(qt)+z(qt)$ for all $t$, it means it is enough to solve the first two equations of system (4): after we know $x(t)$ and $y(t)$, we can immediately compute $z(qt)$ using the equality z(qt)=Nx(qt)y(qt)z(qt)=N-x(qt)-y(qt).
@@
tends to zero as $i$ grows. Now we prove that $\xi_{i+1}<\xi_{i}<1$ for all $i\geq 0$. For a certain $i\geq 0$, suppose that $\xi_{i+1}<\xi_{i}$. Then,
@@
Since, by hypothesis, $\mathcal{R}_{0}<1$, then $\frac{c}{b}>1$, and thus, the result holds for any $q>1$. ∎
@@ 4 Conclusion
In this work, we proposed a quantum SIR model considering the concept of $q$-derivative and a particular rule of Mickens’ method. The equilibrium points were found, and the non-negativity and boundedness of solutions were proved. We derived the exact solution of the quantum SIR model and proved asymptotic stability. We end providing some examples that illustrate the exact solution obtained.
@@
The authors declare no conflicts of interest.